OpenStudy (df001):

If x = Arc sin (-1/2), find angle x.

OpenStudy (bohotness):

Sketch a right triangle, angle A at the left. since cos A = 1/x, label the side adjacent to A with 1, hypotenuse with x. So side opposite A must be √(x² - 1). A = arc cos 1/x, thus sin A = sin(arc cos 1/x) = [√(x² - 1)]/x

OpenStudy (df001):

So in other words it's sin inverse (-1/2)?

OpenStudy (bohotness):

YES

OpenStudy (df001):

thanks buddy